A vertical metal cylinder of radius 2 cm and length 2 m is fixed at the lower end and a load of 100 kg is put on it. Find (a) the stress (b) the strain and (c) the compression of the cylinder. Young modulus of the metal = 2 × 1011 N m-2.

Grade:upto college level

1 Answers

Aditi Chauhan
9 years ago
Sol. . p = stress = mg/A e = strain = p/Y Compression ∆L = eL